# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ Abstract.
Using only 34 published tables, we reconstruct five variables (census block, sex, age, race, and ethnicity) in the confidential 2010 Census person records. Using the 38-bin age variable tabulated at the census block level, at most 20.1% of reconstructed records can differ from their confidential source on even a single value for these five variables. Using only published data, an attacker can verify that all records in 70% of all census blocks (97 million people) are perfectly reconstructed. The tabular publications in Summary File 1 thus have prohibited disclosure risk similar to the unreleased confidential microdata. Reidentification studies confirm that an attacker can, within blocks with perfect reconstruction accuracy, correctly infer the actual census response on race and ethnicity for 3.4 million vulnerable population uniques (persons with nonmodal characteristics) with 95% accuracy, the same precision as the confidential data achieve and far greater than statistical baselines. The flaw in the 2010 Census framework was the assumption that aggregation prevented accurate microdata reconstruction, justifying weaker disclosure limitation methods than were applied to 2010 Census public microdata. The framework used for 2020 Census publications defends against attacks that are based on reconstruction, as we also demonstrate here. Finally, we show that alternatives to the 2020 Census Disclosure Avoidance System with similar accuracy (enhanced swapping) also fail to protect confidentiality, and those that partially defend against reconstruction attacks (incomplete suppression implementations) destroy the primary statutory use case: data for redistricting all legislatures in the country in compliance with the 1965 Voting Rights Act.
[…]
@@
Keywords: statistical disclosure limitation, reconstruction attack, record linkage attack, differential privacy, swapping, suppression
@@ 1. Introduction
Data products from the U.S. Decennial Census of Population and Housing are widely used for policy, research, and community planning including the allocation of approximately \$$\$2.8$ trillion in federal spending to state and local governments, nonprofits, businesses, and households [villa_ross:2023]. In order to support these data uses, tens of billions of statistics are published, predominantly at the most granular level of geographic detail—the census block. With so many statistics published at such a fine geographic detail, an important question arises: “How accurately can a data user reconstruct the underlying confidential record-level data from the published tables?” Working against that accuracy are the confidentiality protections used in the 2010 Census publications. For the 1990, 2000 and 2010 Censuses, aggregation, age coarsening, noise infusion via targeted geographic identifier swapping, and, in 2010, partially synthetic data were used as the statistical disclosure limitation (SDL) framework to protect confidentiality [mckenna2018disclosure]. We study the extent to which these SDL procedures limited the accuracy of reconstructed microdata and impeded reidentification. Hence, the related research question is: “To what extent data did aggregation and record swapping limit the accuracy of reconstruction?”
@@
We reconstruct the underlying person-level records (called microdata) for the features (characteristics) labeled census block, sex, age, race, and ethnicity using only a small subset of publicly released tables. We match these reconstructed records to a low-quality commercial database acquired during the conduct of the 2010 Census containing personal identifiers and to a high-quality personal identifier database constructed from an extract of the 2010 Census data themselves. Our unique contributions to this literature are:
@@
  • the first demonstration supported by a national statistical agency that the reconstruction predicted by [Dinur:Nissim:2003:RIW:773153.773173, ] is feasible at scale using its flagship publication;
@@
  • the complete empirical demonstration that separate, incompatible, confidentiality protection frameworks for tabular and microdata publications fail if the tabular data are too detailed;
@@
  • the first mathematical proof requiring no access to confidential data that a large, identifiable subset of reconstructed records are the exact image of the underlying confidential records for the stated feature set;
@@
  • the first mathematical proof of an upper bound on the percentage of reconstructed records that can differ on no more than a single feature value from their confidential image on the stated feature set;
@@
  • the empirical demonstration that neither aggregation nor collapsing age into narrow bins prevents high-precision reidentification of census respondents from tabular data;
@@
  • the first empirical demonstration that reconstructed microdata succeed in reidentifying vulnerable individuals (those with characteristics that differ from the modal person in the relevant universe) with precision rates much higher than statistical baselines and comparable to the precision rates achieved using the confidential data themselves (vulnerable populations are based on racial and ethnic minorities in this work, but they could be other sensitive characteristics like occupancy-code violations, tribal identities, or same sex partners using other 2010 Census publications);
@@
  • the first research team to place the entire reconstruction workflow in the public domain, permitting others, including other statistical agencies, to assess the risk in the many similar products published by other data stewards;
@@
  • strong demonstration that the differential privacy framework used for the 2020 Census in its May 25, 2023 release defends against this attack at the parameter values used to produce the 2020 Census Demographic and Housing Characteristics File–—successor to the 2010 Census Summary File 1, although there may attacks not yet discovered to which its algorithms remain vulnerable.
This paper also shows how the potential choices for the 2020 Census Disclosure Avoidance System—suppression, enhanced swapping, and differential privacy—addressed the risks exposed by our reconstruction and reidentification studies. We included sufficient detail so that readers could review all data needed to judge this for themselves. It makes the paper longer, but it also shows that protecting vulnerable populations is not a matter of just “turning up the swap rate” or doing some suppression. The data steward must have a workable definition of “vulnerable populations,” which is what the leave-one-out analysis presented in this article provides. Then, the data steward must show that a comprehensive framework designed to protect all vulnerable populations actually works. The 2020 Census differential privacy framework succeeds. The other choices do not.
@@
The 2020 Census Disclosure Avoidance System (DAS) took six years to develop. The portion that produced the data comparable to the 2010 Census tables studied in this paper was finalized in November 2022. Experimentation with alternatives to the 2010 Census SDL framework began in 2016. The decision in 2018 to use a differentially private framework for the DAS was based exclusively on reconstruction results available at that time [abowd2018us]; however, continuous research confirmed that reconstruction risk does imply privacy-violating reidentification risk. Our contributions are timely because traditional disclosure limitation experts continue to dispute the efficacy of reconstruction-based attacks [Muralidhar:2022, muralidhar:domingo-ferrer:2023] using incomplete formulations of the problem, and domain experts continue to assert that the methods are no better than guessing [ruggles:vanriper:2022, francis:2022] or ineffective [kenny:et:al:2021]. Many of these critiques are addressed directly in [jarmin:etal:2023, ] and [garfinkel:2023, ]. However, the analysis of how to properly assess the disclosure risk associated with publishing massive tabulations from a single confidential input continues to focus on methods with the same flaws that our experimental attack exploits [hotzetal:2022]. Every major textbook or review article on SDL [willenborg:deWaal:2000, hundepool:et:al:2012, elliot:domingo-ferrer:2018, duncan:etal:2011:statistical] recommends using distinct methods for tabular and microdata publications. However, the format of the data publication is immaterial because, as we show in this paper, tabulations can be converted into microdata. The use of weaker SDL standards for tabulations as compared to microdata is precisely the flaw that our attack exploits and the recommendation that our research challenges.
@@
Section 2 elaborates the legal, ethical, and statistical confidentiality requirements that the Census Bureau’s disclosure avoidance frameworks are meant to implement. Section 3 lays out the complete schematic workflow of our research, describes all input data sources, and provides a reference table that shows the feature sets (characteristics or variables) of every input and output dataset used in this research. Section 4 details the complete reconstruction methodology and explains how to use the integer programs in the replication archive to perform reconstruction on the same public data. Section 5 describes the algorithms used to match data sets in order to assess reconstruction agreement and reidentification risk. Section 6 assesses the solution variability of our reconstructed microdata. Section 7 demonstrates the strong agreement of our reconstructed microdata with the confidential data. Section 8 assesses the reidentification risk of the reconstructed microdata focusing on the accuracy of inferences for vulnerable populations—those whose characteristics differ from the modal person in the relevant universe. Section 9 demonstrates that the SDL used for the 2010 Census did not meet the Census Bureau’s stated standards for that census. Section 10 demonstrates that the differentially private framework used for the 2020 Census successfully addresses the failures of the methods used in 2010. Section 11 concludes.
@@ 2. Legal, Ethical and Statistical Confidentiality Requirements
It is fundamentally important to address the confidentiality breaches arising from the inconsistency in the SDL methods that were used for past U.S. decennial censuses and many other publications. For the 1990, 2000, and 2010 Censuses, the primary SDL framework was household-level record swapping [mckenna2018disclosure]. Once the record swapping was implemented, there were separate, additional requirements for confidentiality protection of tabular summaries [mckenna2018disclosure] and microdata [mckenna:2019:microdata]. For tabular summaries, publications used the universe of records (no sampling), created tables down to the census block level, and used detailed schemas for demographic variables. Tabular summaries in the 2010 Census imposed no minimum population or household counts on any tables in the main release—the 2010 Summary File 1 (SF1) [sf1:2010:tech]. In contrast, the public-use microdata sample required sampling, minimum population in a geographic area of at least 100,000 persons, and minimum national population in one-way marginals for demographic variables of at least 10,000 persons [mckenna:2019:microdata]. These requirements are inconsistent if the published tabular data can be used to reconstruct an accurate image of the underlying confidential microdata record that includes a geographic identifier with no minimum population, has a record for every person in the census, and includes demographic information for groups with national populations as small as one.
@@
If an accurate reconstruction of the record-level data is possible from tabular summaries, then the rules adopted for the 2010 Census disclosure avoidance, noted in the paragraph above, would require sampling to introduce deliberate sampling error, the suppression of most census block-level data even if the aforementioned 100,000 population threshold were significantly relaxed, and the suppression or aggregation of many race categories even if the population threshold of 10,000 were significantly relaxed. Thus, the application of technology used in the 2010 Census to the 2020 Census data would have resulted in substantial data loss compared to modern methods based on the differential privacy framework.
@@
Furthermore, if an accurate reconstruction of the record-level data is possible from the tabular summaries, it makes the published 2010 Census data susceptible to a reconstruction-abetted reidentification attack in which an attacker reconstructs all or parts of the record-level confidential database from the publicly available information and combines these reconstructed records with an external source of person-level or household-level data containing personal identifiers, thus potentially reidentifying the respondent or another person in the household and learning response data associated with that person. This is a traditional attack vector that has been recognized by statistical agencies [spwp22, mckenna:2019:reid], the National Institute for Standards and Technology [garfinkel:near:etal:2023], and general researchers [rocher:et:al:2019, dick:etal:2023]. When there is sufficient detail in the reconstructed records and there are enough common variables in the reconstructed and external microdata, the attacker may infer previously unknown person or household-level attributes from the reconstructed database with high accuracy, thus associating these characteristics with the individuals or households in their external dataset. Using census data to learn specific responses supplied by identifiable individuals is a prohibited, nonstatistical use as defined in the controlling statutes. [^1] Within the U.S., federal policy on statistical and nonstatistical uses is governed by Statistical Policy Directive 1, now codified in the Confidential Information Protection and Statistical Efficiency Act of 2018 in 44 U.S. Code §§ 3561-4, in particular, the definitions in 3561(8) of “nonstatistical purpose” and (12) “statistical purpose.” For the Census Bureau such prohibited uses are also codified in The Census Act of 1954 (as amended) in 13 U.S. Code §§ 8(b) & 9, See Section 2.1. It is the obligation of statistical agencies to prevent or impede such uses.
@@
In order to understand what we mean by a prohibited, nonstatistical inference, we must also be clear about an allowable scientific, statistical, or generalizable inference. The most straightforward way to do this is by using concepts from robust statistics, specifically leave-one-out (LOO) estimation and inference [wasserman:2010]. Inferences about personal characteristics consist of associating an attribute measured in a survey or census with a particular individual. When such inferences are based on estimators that exclude only the individual under study, they are called LOO inferences. LOO inferences cannot be privacy violations in our analysis because they cannot depend on the particular individual’s confidential data—it was not used in the calculation. This connection between robust inference and privacy analysis was first noticed by [dwork:lei:2009, ], and it is now well-accepted in social science, statistics and computer science. It is a form of causal inference in the sense of [imbens:rubin:2015, ] that has also been applied to confidentiality protection in machine learning [ye:etal:2023].
@@
LOO inference is, by construction, generalizable scientific inference. Non-LOO inference is a privacy violation if it is too precise. The difference in precision between LOO inferences and inferences based on estimators that include the particular individual’s data measures the extent to which the individual’s data caused the inference that associates the feature value with the identifiable person. Therefore, when the difference in precision between the LOO inference and the non-LOO inference is large, there is a strong presumption that the non-LOO inference is a confidentiality violation because the gain in precision from the non-LOO inference is provably due to the presence of data about the individual under study in the estimator used to make the inference. Put directly, the analyst’s precision gain was caused (again, in the sense of [imbens:rubin:2015]) by the use of data supplied by that person. In this paper we evaluate the efficacy of our attack by focusing on situations where the data strongly suggest that inferences based on the published 2010 Census data are much more precise than LOO inferences would be. That is, in the language of causal inference, the observed outcome is a statistic officially published from 2010 Census data and the counterfactual outcome is an approximation of the same statistic published after deleting that individual’s record from the input data.

### v2

# A Simulated Reconstruction and Reidentification Attack on the

@@ Abstract.
For the last half-century, it has been a common and accepted practice for statistical agencies, including the United States Census Bureau, to adopt different strategies to protect the confidentiality of aggregate tabular data products from those used to protect the individual records contained in publicly released microdata products. This strategy was premised on the assumption that the aggregation used to generate tabular data products made the resulting statistics inherently less disclosive than the microdata from which they were tabulated. Consistent with this common assumption, the 2010 Census of Population and Housing in the U.S. used different disclosure limitation rules for its tabular and microdata publications. This paper demonstrates that, in the context of disclosure limitation for the 2010 Census, the assumption that tabular data are inherently less disclosive than their underlying microdata is fundamentally flawed. The 2010 Census published more than 150 billion aggregate statistics in 180 table sets. Most of these tables were published at the most detailed geographic level—individual census blocks, which can have populations as small as one person. Using only 34 of the published table sets, we reconstructed microdata records including five variables (census block, sex, age, race, and ethnicity) from the confidential 2010 Census person records. Using only published data, an attacker using our methods can verify that all records in 70% of all census blocks (97 million people) are perfectly reconstructed. We further confirm, through reidentification studies, that an attacker can, within census blocks with perfect reconstruction accuracy, correctly infer the actual census response on race and ethnicity for 3.4 million vulnerable population uniques (persons with race and ethnicity different from the modal person on the census block) with 95% accuracy. Having shown the vulnerabilities inherent to the disclosure limitation methods used for the 2010 Census, we proceed to demonstrate that the more robust disclosure limitation framework used for the 2020 Census publications defends against attacks that are based on reconstruction. Finally, we show that available alternatives to the 2020 Census Disclosure Avoidance System would either fail to protect confidentiality, or would overly degrade the statistics’ utility for the primary statutory use case: redrawing the boundaries of all of the nation’s legislative and voting districts in compliance with the 1965 Voting Rights Act. You are reading the full technical report. For the summary paper see https://doi.org/10.1162/99608f92.4a1ebf70.
@@
[…]
@@ 1. Introduction
Data products from the U.S. Decennial Census of Population and Housing are widely used for policy, research, and community planning including the allocation of approximately \$$\$2.8$ trillion in federal spending to state and local governments, nonprofits, businesses, and households (villa_ross:2023). In order to support these data uses, tens of billions of statistics are published, predominantly at the most granular level of geographic detail—the census block, a primitive geographic unit with precisely defined geographical boundaries. In 2010, the U.S. was subdivided into 11,078,297 mutually exclusive census blocks that tessellated the country; of these 6,207,027 had positive populations ranging from one person to more than 1,000 persons (see Section 6.2 for detailed statistics). With so many statistics published at such a fine geographic detail, an important question arises: “How accurately can a data user reconstruct the underlying confidential record-level data from the published tables?”
Highly accurate reconstructions of the confidential record-level responses are potential confidentiality breaches because record-level data is easy for external parties to examine for rare or unique records. Thus, record-level data may permit direct, reliable reidentification of census responses, which statutes specifically prohibit the Census Bureau from enabling (see Section 2.2). In addition, many census respondents, if they worry that their confidentiality may be breached, may have meaningful concerns about providing their data. For example, age, sex, race, and ethnicity data about children are often missing in commercial databases (collected and maintained by private companies) due to legal restrictions motivated by the sensitivity of information about children, but the U.S. census collects such data on children and publishes statistical summaries at large scale.
Working against record-level accuracy are the confidentiality protections used in census publications. For the 1990, 2000 and 2010 Censuses, aggregation, age coarsening, noise infusion via targeted geographic identifier swapping, and, in 2010 for group quarters only, partially synthetic data were used as the statistical disclosure limitation (SDL) framework to protect confidentiality (mckenna:2018:disclosure). We study the extent to which these SDL procedures limited the accuracy of reconstructed microdata and impeded reidentification. Hence, the related research question is: “To what extent did data aggregation and record swapping limit the accuracy of reconstruction?”
@@
We reconstruct the underlying person-level records (called microdata) for the features (characteristics or variables) labeled census block, sex, age, race, and ethnicity using only a small subset of publicly released tables. We match these reconstructed records to a lower-quality commercial database acquired during the conduct of the 2010 Census containing personal identifiers and to a high-quality personal identifier database constructed from an extract of the 2010 Census data themselves. Our unique contributions to this literature are:
@@
  • the first demonstration supported by a national statistical agency that the reconstruction predicted by Dinur:Nissim:2003:RIW:773153.773173 is feasible at scale using its flagship publication (see Sections 6.2 and 7);
@@
  • the complete empirical demonstration that separate, incompatible, confidentiality protection frameworks for tabular and microdata publications fail if the tabular data are too detailed (see Section 8);
@@
  • the first mathematical proof requiring no access to confidential data that a large, identifiable subset of reconstructed records are identical to their corresponding confidential records for the specific variables we study (see Sections 4.5 and 6.1);
@@
  • the first mathematical proof of an upper bound on the percentage of reconstructed records that can differ on no more than a single feature value from their confidential image on the stated feature set for all feasible reconstructions (see Sections 4 and 6.1);
@@
  • the empirical demonstration that neither aggregation nor collapsing age into a binned-age schema with narrow bins (single year of age from 0 to 21, then 2-5 year age bins up to 85+) prevent high-precision reidentification of census respondents from tabular data (see Section 7);
@@
  • the first empirical demonstration that reconstructed microdata succeed in reidentifying vulnerable individuals (those with characteristics that differ from the modal person—the person with the most commonly occurring characteristics—in the relevant universe) with precision rates much higher than statistical baselines and comparable to the precision rates achieved using the confidential data themselves (vulnerable populations are based on racial and ethnic minorities in this work, but they could be other sensitive characteristics like occupancy-code violations, tribal identities, or same sex partners using other 2010 Census publications) (see Sections 7.1 and 7.3);
@@
  • the first research team to place the entire reconstruction workflow in the public domain, permitting others, including other statistical agencies, to assess the risk in the many similar products published by other data stewards (see replication archive), the agencies responsible for decisions about data curation and publication;
@@
  • strong demonstration that the differential privacy framework used for the 2020 Census in its May 25, 2023 release defends against this attack at the parameter values used to produce the 2020 Census Demographic and Housing Characteristics File–—successor to the 2010 Census Summary File 1 (the main release of demographic tables in the 2010 Census, especially those with very granular age detail), although there may be attacks not yet discovered to which its algorithms remain vulnerable (see Section 9).
This paper also shows how the potential choices for the 2020 Census Disclosure Avoidance System—suppression, enhanced swapping, and differential privacy—addressed the risks exposed by our reconstruction and reidentification studies. We included sufficient detail so that readers could review all data needed to judge this for themselves, including empirical results that demonstrate the extreme difficulty of guarding against reconstruction-abetted attacks by increasing the swap rate or applying traditional suppression rules. To carefully assess the extent of confidentiality violations, the data steward—the entity responsible for maintaining, safeguarding, and publishing summaries of the collected data—must have a workable definition of “vulnerable populations.” To prevent the extraction and nonstatistical use of personal information statistical agencies must periodically analyze their SDL methods because nonstatistical actors, and especially malicious actors, do not publicly advertise their plans or methods.
This article provides provides a workable definition of “vulnerable populations” by introducing leave-one-out analysis, and uses this concept to assess the effectiveness of the SDL methods used in the 2010 Census in preventing nonstatistical inferences based on the published data, in the context of what we call a reconstruction-abetted reidentification attack. Given a specific attack or set of attacks, the data steward must show that a comprehensive framework designed to protect all vulnerable populations actually works. We first show that the 2010 Census disclosure avoidance system was insufficient to stop the reconstruction-abetted reidentification attack from achieving many egregious confidentiality violations. The risk posed by this attack motivated adoption of a differential privacy framework in the 2020 Census. We therefore also show that this differential privacy framework succeeds in defending against the reconstruction-abetted reidentification attack. We also show that other choices—specifically, the suppression system used in the 1980 Census, and enhanced versions of the swapping systems used in the 1990, 2000, and 2010 Censuses—do not.
To carry out the reconstruction-abetted reidentification attack on the 2010 Census, we first attempt to reconstruct the underlying microdata for the features of census block, sex, age, race, and ethnicity from publicly released tables. We then attempt to match those reconstructed records to (a) records in lower-quality commercial databases purchased during the conduct of the 2010 Census with person and address identifiers (representing attackers with the same quality contemporaneously available information in 2010) and (b) records containing a limited subset of variables from the 2010 Census itself (person identifier, census block, sex and age; representing attackers with higher-quality contemporaneous external information). We used shared variables, called “key variables” in (duncan:etal:2011:statistical, p. 20) or “quasi-identifiers” in (garfinkel:near:etal:2023, p. 49), to determine how accurately we could link records and then infer the race and ethnicity of the persons in (a) the commercial records and (b) the quasi-identifier-only version of the 2010 Census records.
We compare the results in several scenarios against baseline attackers who predicted using either (a) the most common race and ethnicity pair in the block (modal prediction from public tables) or (b) the race and ethnicity pair proportional to the distribution of block-level race by ethnicity combinations (proportional prediction from public tables). In all cases, we assess the prediction accuracy using the full set of 2010 Census variables (person identifier, census block, sex, age, race, and ethnicity). For census block, sex, race, and ethnicity we always used the same schema. For age, we considered two different schemas: the block-level table schema (38 narrow age categories) and the tract-level schema (111 exact age categories). [^1] For the person identifier, we used the same vintage of the Census Bureau’s production record-linkage system to replace the reported name and address with internal identifiers routinely used for person and household record linkage. We note that, in all cases, the attack file does not contain race and ethnicity information, while the reconstructed microdata does not contain person identifiers, so that linking the two files enables the attacker to infer the race and ethnicity of each person to whom a reconstructed record is linked.
The 2020 Census Disclosure Avoidance System (DAS) adopted to guard against reconstruction-abetted reidentification attacks (among others) took six years to develop. The portion that produced the data comparable to the 2010 Census tables studied in this paper was finalized in November 2022. Experimentation with alternatives to the 2010 Census SDL framework began in 2016. The decision in 2018 to use a differentially private framework for the DAS was based exclusively on reconstruction results available at that time (abowd2018us); however, continuous research confirmed that reconstruction risk does imply confidentiality-violating reidentification risk.
@@
Our contributions are timely because traditional disclosure limitation experts continue to dispute the efficacy of reconstruction-based attacks (Muralidhar:2022; muralidhar:domingo-ferrer:2023) using incomplete formulations of the problem, and domain experts continue to assert that the methods are no better than guessing (ruggles:vanriper:2022; francis:2022) or ineffective (kenny:et:al:2021; ruggles2024privacy). Many of these critiques are addressed directly in jarmin:etal:2023 and garfinkel:2023. However, the analysis of how to properly assess the disclosure risk associated with publishing massive tabulations from a single confidential input continues to focus on methods with the same flaws that our experimental attack exploits (hotzetal:2022). Every major textbook or review article on SDL (willenborg:deWaal:2000; hundepool:et:al:2012; elliot:domingo-ferrer:2018; duncan:etal:2011:statistical) recommends using distinct methods for tabular and microdata publications. However, the format of the data publication is immaterial because, as we show in this paper, tabulations can be converted into microdata, and an attacker can, using only publicly available information, derive high-quality bounds on how closely the microdata created in this way resembles the microdata used to calculate the tabulations. The use of weaker SDL standards for tabulations as compared to microdata is precisely the flaw that our attack exploits and the recommendation that our research challenges.
@@
Section 2 elaborates the legal, ethical, and statistical confidentiality requirements that the Census Bureau’s disclosure avoidance frameworks are meant to implement. Section 3 lays out the complete schematic workflow of our research, describes all input data sources, and provides a reference table that shows the feature sets (characteristics or variables) of every input and output dataset used in this research. Section 4 lays out the reconstruction methodology. Section 5 describes the matching and reidentification methodology. Section 6 assesses the solution variability of our reconstructed microdata and demonstrates the strong agreement of our reconstructed microdata with the confidential data. Section 7 assesses the reidentification risk of the reconstructed microdata focusing on the accuracy of inferences for vulnerable populations—those whose characteristics differ from the modal person in the relevant universe, and especially those who are unique on sex and binned age, within their census block. Section 8 demonstrates that the SDL used for the 2010 Census did not meet the Census Bureau’s stated standards for that census. Section 9 demonstrates that the differentially private framework used for the 2020 Census successfully addresses the failures of the methods used in 2010. Section 10 concludes.
@@ 2. Legal, Ethical and Statistical Confidentiality Requirements
2.1. Defining and diagnosing confidentiality violation in 2010 SDL methods
Data from the 2010 Census were published and made available to the public in two primary forms: tabulations (low-dimensional statistical summary statistics, like “Number of people residing in California who are White alone, male, and 26 years old”) and microdata (in a “public-use microdata sample”; individual-level census 51 responses released as part of a “public-use microdata sample). Both tabular summaries and microdata were based on the same Hundred-Percent Detail File (HDF)—microdata created by applying record-level swapping to the collected, edited, and imputed microdata—but the disclosure requirements specified by the Census Bureau in 2010 for tabular and microdata releases were very different. It is fundamentally important to address the confidentiality breaches arising from this inconsistency in the SDL methods that were used for past U.S. decennial censuses. For the 1990, 2000, and 2010 Censuses, the primary SDL framework was household-level record swapping (mckenna:2018:disclosure). After disclosure avoidance based primarily on record swapping was implemented, there were separate, additional requirements for confidentiality protection of tabular summaries (mckenna:2018:disclosure) and microdata (mckenna:2019:microdata). To see the inconsistency most clearly, we quote from the technical documentation of the 2010 Census data products.
For tabular summaries, the Census Bureau required:
“Disclosure avoidance is the process of disguising data to protect confidentiality. A disclosure of data occurs when someone can use published statistical information to identify an individual who provided information under a pledge of confidentiality. Using disclosure avoidance, the Census Bureau modifies or removes all of the characteristics that put confidential information at risk for disclosure. Although it may appear that a table shows information about a specific individual, the Census Bureau has taken steps (such as data swapping) to disguise the original data while making sure the results are useful. …
Data swapping is a method of disclosure avoidance designed to protect confidentiality in tables of frequency data (the number or percentage of the population with certain characteristics). Data swapping is done by editing the source data or exchanging records for a sample of cases. A sample of households is selected and matched on a set of selected key variables with households in neighboring geographic areas (geographic areas with a small population) that have similar characteristics (same number of adults, same number of children, etc.). Because the swap often occurs within a geographic area with a small population, there is no effect on the marginal totals for the geographic area with a small population or for totals that include data from multiple geographic areas with small populations. Because of data swapping, users should not assume that tables with cells having a value of one or two reveal information about specific individuals.” (sf1:2010:tech)
For public-use microdata samples, the Census Bureau required:
“Each microdata file is a 10-percent systematic sample of the full census population. … Using disclosure avoidance, the Census Bureau modifies or removes all of the characteristics that put confidential information at risk for disclosure. Although it may appear that a table shows information about a specific individual, the Census Bureau has taken steps to disguise the original data while making sure the results are useful. The Census Bureau’s internal Disclosure Review Board monitors the disclosure review process and sets the confidentiality rules for all data releases. The main disclosure avoidance method used is to limit the geographic detail shown in the files. A geographic area must have a minimum population of 100,000 to be fully identified. A minimum threshold of 10,000 for the national population (excluding Puerto Rico) was set for identification of groups within categorical variables in the state-level PUMS files. Confidentiality is protected, in part, by the use of the following processes: data-swapping, synthetic data, top-coding of selected variables, age perturbation for large households, and reduced detail on some categorical variables.” (PUMS:2010:tech)
@@ 2.1. Defining and diagnosing confidentiality violation in 2010 SDL methods
For tabular summaries, publications used the universe of records (no sampling), created tables down to the census block level, and used detailed schemas for demographic variables. Tabular summaries in the 2010 Census imposed no minimum population or household counts on any tables in the main release—the 2010 Summary File 1 (SF1) (sf1:2010:tech). In contrast, the public-use microdata sample required sampling, minimum population in a geographic area of at least 100,000 persons, and minimum national population in one-way marginals for demographic variables of at least 10,000 persons (mckenna:2019:microdata). These requirements are inconsistent if the published tabular data can be used to reconstruct an accurate image of the underlying confidential microdata record that includes a geographic identifier with no minimum population, has a record for every person in the census, and includes demographic information for groups with national populations as small as one. We prove that such reconstruction is feasible, and that an attacker can verify the correctness of the reconstruction without access to any confidential data (see Section 4). We further show that 97 million confidential records can be reconstructed with perfect block-level accuracy (see Section 6.2).
@@
If an accurate reconstruction of the record-level data is possible from tabular summaries, then the rules adopted for the 2010 Census disclosure avoidance, noted in the paragraphs above, would require that the same disclosure avoidance procedures applied to the 2010 PUMS also be applied to the 2020 tabular summaries. These include sampling to introduce deliberate sampling error, the suppression of most census block-level data even if the aforementioned 100,000 population threshold were significantly relaxed, and the suppression or aggregation of many race and ethnicity categories even if the population threshold of 10,000 were significantly relaxed. Thus, consistent application of the technology used in the 2010 Census to the 2020 Census data would have resulted in substantial data loss compared to modern methods based on the differential privacy framework.
@@
Furthermore, if an accurate reconstruction of the record-level data is possible from the tabular summaries, it makes the published 2010 Census data susceptible to a reconstruction-abetted reidentification attack in which an attacker reconstructs all or parts of the record-level confidential database from the publicly available information and combines these reconstructed records with an external source of person-level or household-level data containing personal identifiers, thus potentially reidentifying the respondent or another person in the household and learning response data associated with that person. This is a traditional attack vector that has been recognized by statistical agencies (spwp22; mckenna:2019:reid), the National Institute for Standards and Technology (garfinkel:near:etal:2023), and general researchers (rocher:et:al:2019; dick:etal:2023). When there is sufficient detail in the reconstructed records and there are enough common variables in the reconstructed and external microdata, the attacker may infer previously unknown person or household-level attributes from the reconstructed database with high accuracy, thus associating these characteristics with the individuals or households in their external dataset. Using census data to learn specific responses supplied by identifiable individuals is a prohibited, nonstatistical use as defined in the controlling statutes. [^2] Within the U.S., federal policy on statistical and nonstatistical uses is governed by Statistical Policy Directive 1, now codified in the Confidential Information Protection and Statistical Efficiency Act of 2018 in cipsea, in particular, the definitions in § 3561(8) of “nonstatistical purpose” and § 3561(12) “statistical purpose.” For the Census Bureau such prohibited uses are also codified in The Census Act of 1954 (as amended) in title13 (see Section 2.2). It is the obligation of statistical agencies to prevent or impede such uses. We show how to perform these reidentification attacks (see Section 5), and demonstrate that the resulting reidentifications are far more accurate than statistical baselines, confirming that the reidentifications are prohibited nonstatistical uses of the census data (see Section 7).
@@
In order to understand what we mean by a prohibited, nonstatistical inference, we must clearly define an allowable scientific, statistical, or generalizable inference. The most straightforward way to do this is by using concepts from robust statistics, specifically leave-one-out (LOO) estimation and inference (wasserman:2010). Inferences about personal characteristics consist of associating an attribute measured in a survey or census with a particular individual. When such inferences are based on estimators that exclude only the individual under study, they are called LOO inferences. LOO inferences cannot be confidentiality violations in our analysis because they cannot depend on the particular individual’s confidential data—it was not used in the calculation. This connection between robust inference and privacy analysis was first noticed by dwork:lei:2009, where they explored this idea while allowing for attackers to have arbitrary amounts of side information and use arbitrarily sophisticated attacks. Here, we apply the same intuition to non-worst-case settings in which attacker information is limited to specific known data sets, and the attacker uses a specific realizable attack—the reconstruction-abetted reidentification attack—rather than all possible attacks. LOO is a form of causal inference in the sense of imbens:rubin:2015 that has also been applied to confidentiality protection in machine learning (ye:etal:2023). We acknowledge that this approach to defining inferences that are not confidentiality-violating can be counterintuitive and may conflict with the common view that any unwanted or harmful inference about a person based on a publication implies a confidentiality breach. We agree that unwanted and harmful inference LOO inferences are an important matter for policy consideration, but these inferences are neither “identifying” nor confidentiality violating because the target person’s data were not used at all in the publication. We view the selection of which generalizable inferences to permit as a policy issue largely separate from the goal of disclosure avoidance systems and outside the scope of this paper.
@@
LOO inference is, by construction, generalizable scientific inference. Non-LOO inference is a confidentiality violation if it is too precise. The difference in precision between LOO inferences and inferences based on estimators that include the particular individual’s data measures the extent to which the individual’s data caused the inference that associates the feature value with the identifiable person. Therefore, when the difference in precision between the LOO inference and the non-LOO inference is large, there is a strong presumption that the non-LOO inference is a confidentiality violation because the gain in precision from the non-LOO inference is provably due only to the presence of data about the individual under study in the estimator used to make the inference. Put directly, the analyst’s precision gain was caused (again, in the sense of imbens:rubin:2015) by the use of data supplied by that person. In this paper we evaluate the efficacy of our attack by focusing on situations where the data strongly suggest that inferences based on the published 2010 Census data are much more precise than LOO inferences would be. In the language of causal inference, the observed outcome is a statistic officially published from 2010 Census data and the counterfactual outcome is an approximation of the same statistic published after deleting that individual’s record from the input data, and large differences in inference precision between these two statistics are confidentiality breaches.